The mass of DNA collected from the compost soil was 1.523 x10-6g and from the garden soil was 3.767 x10-6g.

Original mass of compost soil = 0.222g

Original mass of garden soil = 0.27g

Assuming:

1) 100% recovery of the cellular DNA from your bacteria; and
2) the mass of DNA in a single bacterium weighs 5.2·10-15 g (true for E. coli at least)

... how many bacteria were present in your soil sample? How many per g of soil? How doesthis compare to what is reported in the scientific literature?
Answer #1

Given that the mass of DNA of single bacteria is 5.2*10-15gram.

Mass of DNA collected from compost soil is 1.523*10-6

So number of bacteria in the compost soil = 1.523*10-6 / 5.2*10-15 = 1.523 * 109/5.2

= 2.9*108.

Mass of DNA collected from the garden is 3.767*10-6

So number of bacteria in the garden soil = 3.767*10-6 / 5.2*10-15 =3.767 * 109/5.2

=7.2*108.

Number of bacteria present in per gram of compost soil = bacteria present in  0.222g / 0.222 = 2.9*108 / 0.222 = 13.19 * 108 = 1.31*109

Number of bacteria present in per gram of graden soil = bacteria present in 0.27g / 0.27 = 7.2*108 / 0.27= 26.6 * 108 = 2.66*109

According to the scientific literature number of bacteria present in the per gram of soil is 109.

So the observed number of bacteria in a per gram of soil is consistent with the scientific literature.

Please write to me if you need more help.

Please hit the thumbs up button